# DRL-Based Resource Allocation for Motion Blur Resistant Federated Self-Supervised Learning in IoV

### v1

@@ VI Simulation Results
In Fig. 5, we compare our algorithm (DRL-BFSSL) with FedCo[34] and FLSimCLR[27], under the same number of iterations for local training. We also conducted experiments on Non-IID datasets. It can be observed that the classification accuracy of all three methods shows an upward trend after training. Among these methods, the FedCo algorithm starts with the lowest accuracy and continues to perform the worst during subsequent training. In contrast, DRL-BFSSL and FLSimCLR exhibit similar accuracy trends, but overall, DRL-BFSSL outperforms FLSimCLR in Top1 accuracy. The training performance on Non-IID datasets was slightly inferior to that on IID datasets but still superior to the FedCo and FLSimCLR algorithms.
The lower performance of FedCo can be attributed to the process described in FedCo[34], where in round $r$, each training vehicle uploads all stored $k$ values (with a batch size set to 512 in the experiment) to BS with a global queue set to 4096 to update the global direction. Updating the queue with $k$ values from different training vehicles compromises the Negative-Negative consistency requirement in MoCo, resulting in lower accuracy for these methods. Conversely, we also implemented the FLSimCLR, which requires a large number of negative samples. FLSimCLR performs worse than DRL-BFSSL primarily due to its dependency on a large number of negative samples to learn effective representations, whereas in our experiments, we used only a few negative samples. While this approach works well in centralized settings, it faces challenges in federated environments where the distribution of negative samples may not be consistent across different clients. The slightly lower performance on Non-IID data can be attributed to the inherent challenges in ensuring uniform representation across heterogeneous data distributions, which affects the overall consistency and learning efficiency in federated settings.
@@
In Fig. 6, we examine scenarios where 5 and 10 training vehicles participate in each round of DRL-BFSSL and scenarios with different local iterations. The red and green lines in Fig. 6(a) represent scenarios with 5 and 10 training vehicles, respectively. It shows an improvement in model accuracy as few vehicles participate. This indicates that as the number of vehicles participating in each round of DRL-BFSSL increases, the number of iterations decreases, indicating fewer communication rounds between training vehicles and the BS. More interactions between vehicles and the BS lead to better model performance. The red and blue lines represent scenarios with 5 training vehicles participating in DRL-BFSSL per round. The red line indicates that training vehicles perform one local iteration, while the blue line indicates two local iterations. It can be observed that the results of two internal iterations surpass those of one iteration, highlighting the advantage of multiple local training iterations even when using the same set of images for local training. We can also observe that when 10 vehicles participate in each round of DRL-BFSSL, the initial accuracy is the lowest, and as the training rounds increase, it gradually becomes comparable to the accuracy achieved with 5 vehicles. This is because initially, the participating vehicles contribute diverse datasets. However, as the number of iterations increases, the newly added vehicles’ datasets become more similar, resulting in reduced datasets diversity. In this scenario, aggregating a small number of models in the initial stages yields better results because these models represent a wider range of datasets diversity.
@@
As shown in Fig. 6(c), we compare the loss curves of the trained models in Fig. 6(a) with the loss function of 5 vehicles after one round of training on the non-IID datasets. It can be observed that the loss function of each experiment exhibits a decreasing trend. Moreover, when 5 vehicles participate in each round of training with each vehicle conducting 2 iterations of local training, the loss decreases the fastest and reaches the minimum value. At the same time, compared to the curves on the IID training set, the loss function of the model trained on the Non-IID datasets exhibits significant fluctuations in the initial stage. However, as the number of episodes increases, the trend of the loss function gradually becomes similar to that on the IID datasets. This phenomenon can be partially attributed to the characteristics of the non-IID datasets. Due to the uneven distribution and diversity of data on the Non-IID datasets, the model faces greater challenges during training. Therefore, the loss function of the model may exhibit significant fluctuations in the initial stage. As training progresses, the model gradually adapts to the data distribution, leading to a reduction in the fluctuation of the loss function, ultimately stabilizing. Additionally, multiple rounds of internal training help the model better learn data features, allowing it to achieve lower loss values in a shorter time.
@@
Figure 8: Comparison between different algorithms
@@
In Fig. 7, we adopt three aggregation methods, including DRL-BFSSL and two baseline algorithms, to aggregate for obtaining a global model with $1/5$ blurred images. Meanwhile, we introduce $3/5$ of blurred images to obtain a global model using the DRL-BFSSL method. From the figure, it can be observed that the loss curves of all four methods exhibit a decreasing trend. Specifically, baseline1 represents the BS uniformly utilizing the average federated aggregation model parameters. Baseline2 represents the BS discarding models trained by blurry image data, and then using average federated to aggregate model parameters. To validate the impact of blurry images on the model, we have set up baseline3, where $3/5$ of the images experience motion blur.
Based on the results of baseline1 and baseline2, it is evident that although baseline1 does not achieve a stable convergence of loss function, compared to baseline2, its loss function reaches a smaller value. In other words, because baseline2 discards some models during aggregation, baseline2 uses the fewest number of models among these methods. This suggests that the number of models participating in aggregation is more critical than the quality of the models. Finally, our proposed aggregation method of DRL-BFSSL not only reduces the fluctuation of the loss function but also converge to a smaller value. Comparing with baseline3, it can be observed that an increasing number of blurry image data is unfavorable for the convergence of the loss function. However, compared to other aggregation methods, the convergence speed remains relatively stable.
@@
In Fig. 8, we compare the performance of SAC, DDPG, TD3, and Random algorithms in solving the optimization problem, covering the optimal results, computation times, and reward functions. From Fig. 8(a), it is evident that the SAC algorithm rapidly reaches convergence and remains stable, while DDPG and TD3 show similar performance, with TD3 slightly outperforming DDPG. Although DDPG and TD3 exhibit some convergence trends, they are highly volatile and less stable. In contrast, the optimal solution obtained by the Random algorithm is significantly higher than that achieved by the SAC algorithm. Overall, the SAC algorithm can obtain better solutions more quickly and stably. This is mainly because SAC introduces entropy rewards to encourage policy exploration, maintaining a certain level of exploration during training.
In Fig. 8(b), we presents a comparison of calculation times. The SAC algorithm reaches a stable value in terms of computation times, while the calculation times of the DDPG algorithm fluctuate significantly. TD3 algorithm shows a similar trend to DDPG algorithm, but with smaller fluctuations. The calculation times for both DDPG algorithm and TD3 algorithm fluctuate around the stable value achieved by SAC. The Random algorithm consistently has higher calculation times than the SAC algorithm, indicating that it has not balanced the optimal problem with calculation times. We aim to have sufficient calculation times to facilitate the training of the global model while minimizing the value of optimization problem. The SAC algorithm excels in this aspect as it can utilize each update more efficiently, enhancing training efficiency. Fig. 8(c) shows the reward functions for each algorithm. The overall reward value of the SAC algorithm is higher than that of DDPG, TD3, and the Random algorithm. The DDPG algorithm might be affected by gradient instability when updating policies, leading to significant fluctuations during its convergence process. In contrast, TD3 algorithm exhibits less fluctuation but still has lower reward values than SAC algorithm. The Random algorithm, due to its inherent randomness, cannot effectively explore the solution space.
@@
In Fig. 9, we compare the performance of the SAC algorithm under different numbers of interfering vehicles, including the value of optimization problem, calculation times and the reward function. Overall, the results of the three can converge to relatively stable values. Specifically, Fig. 9(a) illustrates the optimal solutions under the SAC algorithm in the presence of varying numbers of interfering vehicles. When there are no interfering vehicles, the optimal solution converges to a minimum value and remains stable. However, with the introduction of interfering vehicles, the optimal solution noticeably increases. Despite this increase, as the number of interfering vehicles continues to grow, the optimal solution does not vary significantly, but ultimately converging to a relatively small and stable value. This behavior can be attributed to the SAC algorithm’s ability to adapt to increased complexity and interference, effectively maintaining a robust performance even under more challenging conditions. Fig. 9(b) shows that when there are no interfering vehicles, the calculation times converge to its minimum value the fastest. As the number of interfering vehicles increases, the convergence value also rises. This is because when the number of interfering vehicles increases, the system needs to allocate more energy to the vehicles for training. Fig. 9(c) shows a comparison of the reward function throughout the process. In general, when there are no interfering vehicles, the reward function can quickly converge to a higher value. However, as the number of interfering vehicles increases, the reward function decreases. At the same time, it can be seen that as long as interfering vehicles are introduced, the overall system reward will significantly decrease. This indicates that once interfering vehicles are added, the overall system reward will drop significantly. The increase in interfering vehicles makes management and energy allocation more complex, thereby reducing the efficiency and effectiveness of the SAC algorithm.
@@
With the absence of interfering vehicles, we train models using both the SAC, DDPG, TD3 and Random algorithms and saved their parameters for testing. As shown in Fig. 10, the overall results indicate that the optimization problem solutions produced by the SAC algorithm are smaller and more stable compared to those produced by the other algorithms. This suggests that the SAC algorithm outperforms the other algorithms in terms of both stability and performance. This superiority is likely due to the entropy regularization strategy employed by the SAC algorithm, which allows for a better balance between exploration and exploitation, leading to more stable convergence to superior solutions. In contrast, the DDPG and TD3 algorithm may be more prone to getting stuck in local optima during policy optimization, resulting in lower quality and less stable solutions.
@@ VII Conclusions
In this paper, we proposed the DRL-BFSSL algorithm, which ensures privacy protection and avoid the data labeling. It also addressed aggregation challenges causing by motion blur in IoV. Additionally, we introduced a resource allocation strategy based on the SAC algorithm to minimize energy consumption and delay for computing and model transmission, during this process, employed packet error rate to set a minimum transmission power to ensure that the BS can successfully receive the local models uploaded by vehicles. The conclusions are summarized as follows:
  • We proposed the DRL-BFSSL algorithm, which protects user privacy, eliminates the need for labels in local training and achieves higher classification accuracy on both IID and Non-IID datasets.

### v2

@@ VI-B Testing settings
TABLE II: Parameters for PSO algorithm
Parameter Value
Maximum iterations 100
Inertia weight 0.2
Personal learning coefficient 0.1
Social learning coefficient 0.1
Positions bounds $[0.0001,1]$
VI-C Comparative algorithms Analysis
VI-C1 Particle Swarm Optimization (PSO)
A population-based optimization algorithm that simulates the behavior of bird flocks or fish schools in search of food. Each particle moves in the solution space using position and velocity, where position represents a solution and velocity determines the direction of movement [67]. And the detailed parameter settings are shown in Table II.
VI-C2 Heuristic Gibbs Greedy Algorithm (HGGA)
An optimization algorithm that combines heuristic methods and greedy strategies, primarily used for combinatorial optimization problems. It generates candidate solutions through heuristic rules and Gibbs sampling, prioritizing decisions that significantly impact solution quality to accelerate the search process [68].
VI-C3 Hierarchical Genetic Recombination Algorithm (HGRA)
An optimization algorithm that combines heuristic methods and random search, primarily used for solving complex combinatorial optimization problems. It introduces randomness and heuristic rules during the search process to avoid getting stuck in local optima and improve global search capability [69].
VI-C4 DDPG
An DRL algorithm suitable for continuous action spaces, combining the concepts of DQN and the Actor-Critic architecture. It employs an Actor-Critic structure where the Actor selects actions and the Critic evaluates their values. DDPG updates the policy deterministically and uses experience replay and target networks to stabilize the training process.
VI-C5 Twin Delayed Deep Deterministic Policy Gradient (TD3)
An improved version of DDPG designed to address the overestimation bias present in DDPG. It enhances stability and performance by introducing twin Critic networks, target policy smoothing, and delayed updates. TD3 reduces noise interference during policy updates, demonstrating better convergence and robustness, particularly for complex continuous control tasks.
VI-C6 Proximal Policy Optimization (PPO)
A policy optimization algorithm that is a simplified version of Trust Region Policy Optimization (TRPO), balancing performance and stability. It stabilizes the training process by clipping policy updates, limiting the change between old and new policies. PPO is applicable to both discrete and continuous action spaces, and its simplicity and superior performance have led to its widespread use in various reinforcement learning tasks.
VI-D Complexity analysis
Regarding the complexity analysis of the DRL-BFSSL algorithm, which combines SAC and SSL, the computational complexity mainly arises during model training.
VI-D1 Complexity analysis of SAC
SAC includes both a policy network (actor) and two value networks (critic). Each network has $L$ layers (4 hidden layers and 2 output layers), and the number of parameters in each layer is $p_{i}$. The computational complexity for a single forward and backward pass is $\mathcal{O}(\sum_{i=1}^{L}p_{i})$. Let $\mathcal{M}$ be the batch size and $K$ be the number of multi-step updates. The computational complexity considering the batch size and multi-step updates is $\mathcal{O}(\mathcal{M}K\sum_{i=1}^{L}p_{i})$.
VI-D2 Complexity analysis of SSL
$\mathcal{M}$ be the batch size, and the number of parameters in each layer is $p_{j}$. The computational complexity of SSL is $\mathcal{O}(\mathcal{M}^{2}p_{j})$.
VI-D3 Complexity analysis of DRL-BFSSL
Since DRL-BFSSL algorithm separates SSL training and resource allocation into two parallel-running components, the overall computational complexity of the proposed algorithm can be expressed as $\mathcal{O}[\text{max}(\mathcal{M}K\sum_{i=1}^{L}p_{i},\mathcal{M}^{2}p_{j})]$.
VI-E Simulation results
In Fig. 5, we compared our proposed DRL-BFSSL algorithm with FedCo [31] and FLSimCLR [23], under the same local iterations. We also conducted simulations on Non-IID datasets based on our proposed DRL-BFSSL algorithm. It can be seen that our proposed DRL-BFSSL algorithm achieves the highest classification accuracy, followed by FLSimCLR, with FedCo having the lowest classification accuracy. At the same time, although the classification accuracy of DRL-BFSSL on Non-IID data is lower than on IID datasets, it still remains higher than that of FedCo.
The lower classification accuracy of FedCo can be attributed to its method described in FedCo, where in each round $r$, each training vehicle uploads all stored $k$ values to BS to form a new global queue for updating the global dictionary. However, updating $k$ values from different training vehicles disrupts the Negative-Negative consistency requirement in MoCo, leading to reduced accuracy for these methods. The classification accuracy of the FLSimCLR algorithm is lower than that of our proposed algorithm, primarily because FLSimCLR, which is based on SimCLR, relies heavily on a large number of negative samples during training. However, in our simulation, given the limited number of images, which did not provide an adequate number of negative samples. Consequently, the classification accuracy of FLSimCLR was lower. Additionally, the slightly lower performance of DRL-BFSSL on Non-IID data compared to IID datasets is primarily due to the inherent challenges in handling heterogeneous data, which affect overall consistency and learning efficiency in federated settings. From another perspective, the stable performance of DRL-BFSSL on Non-IID datasets also indicates that our proposed algorithm has good adaptability and robustness in the face of data heterogeneity that are more common in practical applications.
@@ VI-E Simulation results
In Fig. 6, we compared the Top1 and Top5 classification accuracy and the corresponding trends in the loss function under varying numbers of vehicles and local iterations per round. Comparing the red and green lines in Fig. 6(a), it shows that model accuracy improves as fewer vehicles participate. This indicates that, since the total number of training vehicles is fixed for each training episode, a decrease in the number of vehicles participating in each round of DRL-BFSSL leads to an increase in the number of communication rounds with BS. This, in turn, results in more communication rounds between the training vehicles and BS, and more communication rounds between vehicles and BS lead to better model performance. The red and blue lines represent scenarios with same training vehicles participating in DRL-BFSSL per round. The red line indicates that training vehicles perform one local iteration, while the blue line indicates two local iterations. It can be observed that the results of two local iterations surpass those of one iteration. This indicates that increasing the number of local iterations leads to higher classification accuracy. We can also observe that when 10 vehicles participate in each round of DRL-BFSSL, the initial accuracy is the lowest, and as the training rounds increase, it gradually becomes comparable to the accuracy achieved with 5 vehicles. In the early stages of training, the local models may not be fully trained, their performances are not stable. At this point, aggregating more models could combine the errors of these inadequately trained models into the global model, leading to poor performance. However, as the number of episodes increases and more data is used, the quality of model training stabilizes, and the impact of data heterogeneity on model performance becomes less significant.
@@
Figure 8: Average utility between SAC and other improved traditional algorithms
As shown in Fig. 6(c), we compared the loss curves of the trained models in Fig. 6(a) on IID and Non-IID datasets. It can be observed that the loss function of each simulation exhibits a decreasing trend. Moreover, when 5 vehicles participate in each round of training with each vehicle conducting two local iterations, the loss decreases the fastest. That’s because multiple local iterations help the model better learn data features, allowing it to achieve lower loss values in a shorter time. At the same time, compared to the simulations on IID datasets, the loss function of the model trained on Non-IID datasets exhibits significant fluctuations in the initial stage. However, as the number of episodes increases, the trend of the loss function gradually becomes similar to that on IID datasets. This phenomenon can be partially attributed to the characteristics of Non-IID datasets. Due to the uneven distribution and heterogeneity of data on Non-IID datasets, the model faces greater challenges in initial stage of training. As training progresses, the model gradually adapts to the data heterogeneity, leading to a reduction in the fluctuation of the loss function. From the perspective of curve variability, the loss function curve exhibits certain fluctuations. This phenomenon is due to models being trained across multiple decentralized vehicles, which can introduce variability due to differences in local data heterogeneity, training conditions, and the frequency of model updates.
@@
Figure 9: Comparison between different RL algorithms
@@
[…]
Increasing the number of local calculation times typically enhances classification accuracy; in other words, higher CPU frequencies often significantly improve classification performance. Therefore, in Fig. 8, we presented a comparative result of the utility between the SAC algorithm and traditional optimization methods based on heuristic and evolutionary algorithms, namely PSO, HGGA, and HGRA. To maximize resource utilization, lower average utility values reflect the superiority of an algorithm in resource allocation. As shown in the figure, SAC achieves the lowest utility within a reasonable range of CPU frequency selection. This indicates that SAC effectively balances computing power and energy consumption when optimizing resource allocation, thereby achieving higher classification performance. This advantage stems from its use of DRL strategies, which allow the algorithm more flexibly to the varying computational demands and resource constraints in dynamic environments, thereby enhancing overall performance. Additionally, in the inset of Fig. 8, we zoomed in the interval $[2850,3100]$, which clearly illustrates that during the first 1000 episodes, the SAC algorithm rapidly reaches a local minimum utility, followed by a gradual increase and eventual stabilization. This shows that SAC quickly finds the highest CPU frequency within a reasonable range during the initial phase. However, as training progresses, the algorithm gradually discovers a balance point between computational performance and energy consumption, leading to a slight decrease in CPU frequency and eventual stabilization at this compromise point.
@@
In Fig. 9, we compared the performance of SAC, DDPG, TD3 and PPO algorithms in solving the optimization problem, covering the value of optimization problem, calculation times, and reward functions. From Fig. 9(a), it is evident that the SAC algorithm rapidly reaches convergence and remains stable. DDPG and TD3 show similar performance, but TD3 slightly outperforming DDPG. The PPO algorithm exhibits a fluctuation level similar to TD3, but it fails to explore a suitable value of optimization problem. Overall, the SAC algorithm can obtain better solutions more quickly and more stably compared to other methods. This is mainly because SAC introduces entropy rewards to encourage policy exploration, maintaining a certain level of exploration during training. In Fig. 9(b), we presented a comparison of calculation times. The calculation times obtained by PPO algorithm are high. The calculation times for the SAC, DDPG, and TD3 algorithms fluctuate around a certain value, with SAC having the smallest fluctuation, followed by TD3, and then DDPG. Fig. 9(c) shows the reward functions for each algorithm. The reward function of the SAC algorithm converges quickly and tends to be stable. The TD3 algorithm ultimately achieves a reward value that is second only to the SAC algorithm but exhibits some minor fluctuations. In contrast, the DDPG algorithm shows larger fluctuations in reward values and does not display a clear convergence trend. The PPO algorithm achieves relatively stable reward values, but its rewards are significantly lower than those of the SAC, DDPG, and TD3 algorithms. The SAC algorithm uses an entropy regularization strategy that better balances exploration and exploitation, allowing it to find the optimal solution more quickly and stably. The TD3 algorithm introduces instability in some scenarios through its dual policy delay update mechanism. Meanwhile, the DDPG algorithm is sensitive to hyper-parameters and lacks sufficient exploration mechanisms, making it prone to getting stuck in local optima and difficult to converge. PPO’s policy updates are more conservative, and while it performs well in discrete action spaces, it does not achieve optimal performance in complex continuous action spaces. We aim to have sufficient calculation times to facilitate the training of the global model while minimizing the value of optimization problem. From Fig. 9(b), we can see that the SAC algorithm does not achieve the highest calculation times. However, as shown in Fig. 9(c), the final reward value of the SAC algorithm is the highest among the four methods. This indicates that the SAC algorithm does not focus solely on maximizing the calculation times during training. Instead, it aims to achieve a relatively reasonable calculation times and pursue higher rewards.
In Fig. 10, we compared the performance of SAC algorithm under different numbers of interfering vehicles, including the value of optimization problem, calculation times and the reward function. Overall, the results of the three can converge to relatively stable values.
@@
Specifically, Fig. 10(a) and Fig. 10(b) illustrate the value of optimization problem and calculation times under SAC algorithm in the presence of varying numbers of interfering vehicles. When there are no interfering vehicles, the value of optimization problem and the calculation times converge to a minimum value and remains stable quickly. However, with the introduction of interfering vehicles, the value of optimization problem and calculation times noticeably increases, and gradually converging to a stable value. This is because when the number of interfering vehicles increases, the system needs to allocate more energy to the vehicles for training. This behavior of converging can be attributed to the SAC algorithm’s ability to adapt to increased complexity and interference, effectively maintaining a robust performance even under more challenging conditions. Fig. 10(c) shows a comparison of the reward function throughout the process. In general, when there are no interfering vehicles, the reward function can quickly converge to a higher value. However, as the number of interfering vehicles increases, the reward function decreases. At the same time, it can be seen that as long as interfering vehicles are introduced, the overall system reward will significantly decrease. The increase in interfering vehicles makes management and energy allocation more complex, thereby reducing the efficiency and effectiveness of the SAC algorithm.
@@
With the absence of interfering vehicles, we trained models using both the SAC, DDPG, TD3 and PPO algorithms and saved their parameters for testing. We conducted 50 tests and presented the results in Fig. 11. As shown in Fig. 11, the overall results indicate that the value of optimization problem produced by the SAC algorithm are smaller and more stable compared to those produced by the other algorithms. This suggests that the SAC algorithm outperforms the other algorithms in terms of both stability and performance. This superiority is likely due to the entropy regularization strategy employed by the SAC algorithm, which allows for a better balance between exploration and exploitation, leading to more stable convergence to superior solutions. In contrast, the DDPG and TD3 algorithm may be more prone to getting stuck in local optima during policy optimization, resulting in lower quality and less stable solutions.
@@ VII Conclusions
In this paper, we proposed DRL-BFSSL algorithm effectively mitigates motion blur while achieving high classification accuracy under privacy protection and label-free conditions. Additionally, the proposed resource allocation strategy based on the SAC algorithm minimizes system energy consumption and delay. Therefore, the method can be extended to various application fields, including traffic management and environmental monitoring in smart cities, resource allocation and scheduling in industrial automation, and enhancing system safety and stability in autonomous driving. The contributions of this paper can be summarized as follows:
  • We proposed the DRL-BFSSL algorithm, which protects vehicle privacy, eliminates the need for labels in local training and achieves higher classification accuracy on both IID and Non-IID datasets after aggregating.
@@
Although the DRL-BFSSL algorithm has demonstrated good performance in smaller-scale IoV environments, it may face scalability issues in larger or more complex networks. Additionally, this study only estimates motion blur from the perspective of vehicle velocity, and challenges remain in addressing sensitivity to motion blur caused by multiple factors, such as moving environments, whether as well as vehicle change directions etc. In the future, we will further investigate these two aspects to enhance the scalability of the algorithm and the accuracy of motion blur estimation.